# Projecting Assumptions: The Duality Between Sparse Autoencoders and Concept Geometry

### v1

@@ 4.1 SpaDE, or How to Design A Geometry-Driven SAE
We now illustrate, by example, how one could exploit the duality between concept structure and SAE architecture to design novel SAEs tailored to specific use cases. We use the data properties—nonlinear separability and concept heterogeneity—and through the duality, construct one set of sufficient conditions on the SAE to capture both properties, resulting in a novel SAE called SpaDE (Sparsemax Distance Encoder).
  • Nonlinear separability. Nonlinear separability can be captured by SAEs with flexible receptive fields, i.e., latents whose receptive fields can morph into the structure of concepts in data. Introducing competition in the encoder nonlinearity—which manifests as equality constraints on the projection set $\mathcal{S}$ in projection nonlinearities (Sec. 3)—leads to greater receptive field flexibility, since receptive fields for each latent are then defined relative to all other latents. Replacing the linear transformation in projection encoders with Euclidean distances to a set of prototypes in input space further improves flexibility, since it allows receptive fields to better exploit magnitude in their structure (linear transformations place heavy emphasis on direction).
  • Concept heterogeneity. For concept heterogeneity, SAEs must demonstrate adaptive sparsity in their latent representations, i.e., different concepts must be able to activate different numbers of latents. For projection nonlinearities, this implies that the projection set $\mathcal{S}$ must admit points with varying levels of sparsity.
@@
Figure 4: SpaDE projects onto the probability simplex using sparsemax. In this illustrative $3D$ figure, note $\|\bm{x}\|_{0}=3$ for points on the face, $\|\bm{x}\|_{0}=2$ for points on edges lying along subspaces, and $\|\bm{x}\|_{0}=1$ for corners that lie on coordinate axes, demonstrating adaptive sparsity.
@@
To satisfy the desiderata above, we use a simple first-order equality constraint on the projection set $\mathcal{S}$, resulting in the probability simplex $\mathcal{S}=\Delta^{s}=\{\bm{x}\in\mathbb{R}^{s}:\sum_{i}x_{i}=1,\bm{x}\geq\bm{0}\}$. The non-negativity is necessary to explain away data as a combination of features with positive contributions. Projection onto the simplex (see Fig. 4) results in the sparsemax nonlinearity (Martins & Astudillo (2016)):
$\begin{split}\text{Sparsemax}(\bm{v})&=\operatorname*{arg\,min}_{\bm{\pi}\in\Delta^{s}}\|\bm{\pi}-\bm{v}\|_{2}^{2}\\ &=\bm{\Pi}_{\mathcal{S}}\left\{\bm{v}\right\},\;\mathcal{S}=\Delta^{s}.\end{split}$
@@
The probability simplex $\Delta^{s}$ admits representations with any (non-zero) level of sparsity: for any choice of $k\in\{1,2,...,s\}$, there are ($s\choose k$ choices of indices $\mathcal{M}_{k}$ for a $k$-sparse representation, and points $\{\bm{x}\in\mathbb{R}^{s}:x_{i}=0,i\notin\mathcal{M}_{k},\sum_{j\in\mathcal{M}_{k}}x_{j}=1,x_{j}\geq 0\}\subseteq\Delta^{s}$ which admit this level of sparsity, thereby capturing concept heterogeneity. Combining Sparsemax with euclidean distances then yields SpaDE:
$\begin{split}\bm{z}=\bm{f}(\bm{x})&=\text{Sparsemax}(-\lambda d(\bm{x},\bm{W})),\\ \text{where }d(\bm{x},\bm{W}))_{i}&=\|\bm{x}-\bm{W}_{i}\|_{2}^{2}.\end{split}$ (5)
@@
In the above, $\lambda$ is a scaling parameter (akin to inverse temperature), $\bm{W}_{i}$ is the $i^{th}$ column of the encoder matrix $\bm{W}$ which behaves as a prototype (or landmark) in input space since we compute euclidean distance from input $\bm{x}$ to $\bm{W}_{i}$. App. D.4, D.2.3 describe the receptive fields of SpaDE in further detail and show how it captures nonlinear separability and concept heterogeneity. It is also worth noting that although SpaDE appears to be different from traditional SAEs using linear transformations, it subsumes a ReLU SAE via an input-adaptive linear transformation and bias (see App. D.4):
$\text{Sparsemax}(-\lambda d(\bm{x},\bm{W})))$ $=ReLU(\bm{A}(\bm{x})\bm{x}+\bm{b}(\bm{x})),$
where $\bm{A},\bm{b}$ are piecewise constant functions of $\bm{x}$. We note this result is exact (i.e., not a mere Taylor approximation). However, SpaDE differs from a ReLU SAE by using linear transformations defined with respect to a local origin, which is uniquely determined by the set of active SAE latents (App. D.4), similar to recent work on steering (Wu et al., 2025b).
@@
We also note that the outer optimization for SpaDE is K-Deep Simplex (KDS, Tasissa et al. (2023)), a modified dictionary learning technique which incorporates locality into sparse representations. The regularizer from KDS is a distance-weighted $\ell_{1}$ regularizer $\mathcal{R}(\bm{z})=\sum_{i}z_{i}\|\bm{x}-\bm{W}_{i}\|_{2}^{2}$, which encourages prototypes to move closer to data when they are active, increasing sparsity of representation  [^1] This regularizer encourages dictionary atoms to “stick” to the data, addressing a concern raised in recent works Fel et al. (2025); Paulo & Belrose (2025) that the directions learned by SAEs may be out-of-distribution (OOD), potentially contributing to their instability.. The inner optimization for SpaDE is a one-sided sparsity-regularized optimal transport (see App. D.4). The implicit data assumption made by SpaDE is distance-separation of concepts.
@@ 5 Results: Empirical Validation of SAE behavior
We perform a suite of experiments which involve training ReLU, JumpReLU, TopK and SpaDE SAEs on datasets ranging from purely synthetic Gaussian clusters to semi-synthetic formal-language model activations and natural vision model activations. Our synthetic experiments morph data properties from Sec. 4 into simple data structures, allowing us to validate our claims about implicit assumptions in SAEs. Experiments on more naturalistic data seek to demonstrate our claims extend to realistic data settings. We note that we report only a subset of results to emphasize our main claims, but a substantially more exhaustive list of results can be found in App. E.
Figure 5: Effect of Nonlinear Separability on SAEs. Each column represents a different SAE. a) $F_{1}$ scores of the top 5 most monosemantic latents (highest F1 scores), where shaded region is $\pm$1SD, of each SAE on two concepts—orange (linearly separable) and purple (non-linearly separable). SAEs that assume linear separability struggle to capture the nonlinearly separable concept. b) Receptive fields of the most monosemantic latent for each SAE, illustrating how some architectures fail to isolate the nonlinear concept cleanly.Intensity of color indicates strength of SAE latent activation. (c) Matrix of pairwise cosine similarities between sparse codes of different datapoints, and data clusters obtained through spectral clustering on this matrix. SpaDE is able to maintain clear concept boundaries and doesn’t mix distinct features.
@@ 5.1 Separability Experiment
Dataset: We constructed a 2-dimensional dataset with Gaussian clusters (abstraction of concepts) of different magnitudes in order to demonstrate nonlinear separability of concepts in a simple setting which facilitates visualization. The concepts with smaller norm are not linearly separable, while those with larger norm are linearly separable.
Experiments: We train the four SAEs—ReLU, JumpReLU, TopK and SpaDE—on this dataset with a range of sparsity levels. Following our arguments about implicit assumptions in SAEs, we hypothesize that ReLU and JumpReLU will be unable to capture the nonlinearly separable concepts with monosemantic latents.
Observations: Fig. 5 shows how different SAEs fare on this experiment. Row (a) shows top-5 F1 scores of the latents activating for different concepts for each SAE, for varying levels of sparsity. We note that F1 scores roughly increase with increasing $L_{0}$ across models. ReLU and JumpReLU achieve an F1 score of 1 for the separable concept (orange), while their F1 scores are much lower and bounded above (by  0.5) for the nonlinearly separable concept (purple). This is clearly seen in the receptive fields in Row (b), where both these models show half-space receptive fields, which overlap with other concepts in the nonlinearly separable case. TopK performs somewhat poorly on both concepts, showing comparable F1 scores in both cases. This may be due to variability across latents, since the most monosemantic latent (highest F1 score) does capture the separable concept quite well (see the receptive field of TopK). While the receptive fields of TopK are more flexible in shape than ReLU, JumpReLU, they struggle to capture the structure of the nonlinearly separable concept. SpaDE shows a top F1 score of 1.0 for both concepts (perfect precision and recall), with its receptive fields demonstrating locality and the ability to capture concept structure, irrespective of whether or not the concept is linearly separable.
Row (c) in Fig.5 shows the similarity between concepts using the complete sparse code, computed as the cosine similarity of the sparse code between different examples. ReLU and JumpReLU show significant cross-concept correlations, observed as a large block of correlations between concepts $C_{1},C_{2},C_{3}$. Spectral clustering on this matrix (also in row (c)) reveals this behavior—these three concepts are assigned the same color. TopK does marginally better, showing correlations between pairs of concepts. SpaDE shows clear delineation of different concepts with near-perfect block structure and clearly separated concepts in spectral clusters. Note that SpaDE may overspecialize and lead to further subclusters, as seen by two colors within concept 1 in row (c).
@@ 5.2 Heterogeneity Experiment
Figure 6: Effect of Concept Heterogeneity on SAEs. a) Per-concept sparsity as a function of intrinsic dimension. Ideally, an SAE should allocate more latents to higher-dimensional concepts. While some SAEs adapt their sparsity levels, others (e.g., TopK) enforce a fixed sparsity, leading to poor adaptation. Colors indicate per-concept MSE—higher errors (red/yellow) show when an SAE fails to capture a concept effectively. b) Normalized MSE vs. per-concept sparsity. A well-performing SAE should maintain low error across all concepts. TopK SAE only achieves good reconstruction (below the dashed 20% error threshold) when sparsity (fixed for a given model) exceeds intrinsic dimensionality, highlighting its lack of flexibility. c) Cosine similarity between SAE latents, visualized for two sparsity levels. Ideally, latents should cluster cleanly by concept. Some SAEs show cross-concept mixing (off-diagonal similarities), indicating that they fail to learn disentangled representations.
@@
Dataset: We generated Gaussian clusters (again an abstraction for concepts) in a 128-dimensional spaces. The five concepts are heterogeneous—they belong to subspaces with different intrinsic dimensions (6, 14, 30, 62, 126), but they are designed to have isotropic structure within each cluster, and similar total variances across clusters.
Experiments: We trained ReLU, JumpReLU, TopK SAEs and SpaDE on this data with varying sparsity levels. We hypothesize that TopK will not be able to adapt its representations to the intrinsic dimension of each cluster.
Observations: Fig. 6 shows the results of all SAEs on this experiment. In Row (a), TopK shows the same level of sparsity per concept for all concepts (solid lines connect the same model across concepts, while different lines represent models trained with different hyperparameters), irrespective of the intrinsic dimension of the concepts. As intrinsic dimension increases, TopK shows worse reconstruction error per concept, which is color coded in the same plot. In contrast, other SAEs—ReLU, JumpReLU and SpaDE show adaptive sparsity to different extents by adjusting their representations to the intrinsic dimension of each concept. Note that SpaDE can capture the intrinsic dimension nearly perfectly (along the dashed $y=x$ curve) for a specific choice of hyperparameters.
@@
Figure 7: Investigating SAE properties on GPT for formal languages. (a) 3D PCA of model activations and SAE encoder weights, where datapoints are colored by part-of-speech (PoS). Encoder weights are indicated by points for SpaDE and arrows for the other SAEs. (b) Matrix of cosine similarities between pairs of data and pairs of latents (in order) for each SAE. White lines separate different PoS. (c) MSE normalized by PoS variance as a function of sparsity, for each PoS. Inset: cumulative sum of variance (eigenvalues of data correlations) of each PoS, where the effective dimension (variance $>99\%$) of each PoS is shown. (d) Top-20 $F_{1}$-scores for different PoS from each SAE’s latents (a measure of monosemanticity).
@@
Row (b) shows reconstruction error per concept normalized by the variance of that concept. Note that a naïve estimator which predicts the mean of each concept will achieve a normalized MSE of 1. For TopK, normalized MSE goes below $20\%$ (i.e., explains $80\%$ of the variance) for each concept only when $k$ exceeds the dimension of that concept. For example, $d=6$ goes below the dashed line only after $k=8$, similarly for other concepts. Other SAEs are able to stay below the $20\%$ threshold for nearly all concepts across hyperparameters.
@@
Row (c) shows co-occurrence of latents, computed as the cosine similarity between a pair of latents across many data points (of all concepts), for two sparsity levels—a sparse vs. dense one. Each latent is assigned a concept which it activates maximally for (averaged over datapoints of that concept). We note that different concepts use different numbers of latents in ReLU, JumpReLU, and SpaDE. However, there are correlations across concepts in ReLU and JumpReLU (for the dense case), indicating co-occurrence of latents across concepts, which does not occur in the sparse case. Correlations are absent in SpaDE under both cases. TopK uses similar number of latents across concepts, inline with its lack of adaptivity.
@@ 5.3 Formal Languages
Dataset: Building on recent work using formal languages for making predictive claims about language models (Jain et al., 2023; Lubana et al., 2024; Allen-Zhu & Li, 2023), we use this setting as a semi-synthetic setup for corroborating our claims. Specifically, we analyze the English PCFG with subject-verb-object sentence order proposed in Menon et al. (2024).
Experiment: We train 2-layer Transformers (Andrej Karpathy, 2023) from scratch on strings of length 128 tokens from the formal grammar above. SAEs are then trained on activations retrieved from the middle residual stream of the model.
Figure 8: SAE properties on DINOv2 activations. (a) 3-D PCA of data (model activations) colored by class, and SAE encoder weights (points for SpaDE, arrows for other SAEs). (b) Cosine similarities between pairs of data and pairs of SAE latents (in order) for each SAE. White lines separate classes. (c) $F_{1}$ scores of top-5 most monosemantic latents for each SAE, for different classes. Color indicates the class.
Observations: Results are shown in Fig. 7. We first note that different parts of speech (PoS), the core concepts of the grammar, seem to form clusters in a three-dimensional PCA of their representations (see row (a)). SpaDE learns to tile the PoS clusters well. While all SAEs do a good job at making their latents uncorrelated across PoS (first column for each SAE in row (b)), there are co-occurring latents across PoS in all SAEs except SpaDE (cross-concept correlations seen in second column of each SAE in row (b)). PoS seem to have different intrinsic dimensions (number of dimensions to capture $99\%$ of total variance in data, inset in row (c)), which leads to TopK requiring different values of K to explain the data (crosses the $5\%$ normalized mse with differing values of k, row (c)). PoS also appear to have differing levels of linear separability, as ReLU and JumpReLU show lower F1 scores which peak at different levels of sparsity for each concept (row (d)), while SpaDE shows a perfect F1 score of 1 in its most monosemantic latents.

### v2

@@ 4.1 SpaDE, or How to Design A Geometry-Driven SAE
We now use the data properties studied above—nonlinear separability and concept heterogeneity—and through the duality, construct one set of sufficient conditions on the SAE to capture both properties, resulting in a novel SAE called SpaDE (Sparsemax Distance Encoder). See App. D.4 for details. We introduce SpaDE as a geometry-driven SAE to validate our claims about the duality between concept geometry and SAE architecture. Hence, SpaDE is expected to capture concepts better than other SAEs when its data assumptions are met.
Nonlinear separability can be captured by SAE encoders with a competitive projection nonlinearity (allowing flexible receptive fields, shaped by locations of all weights) and compute Euclidean distances to a set of prototypes instead of linear transforms (to better exploit magnitude). For concept heterogeneity, SAEs must demonstrate adaptive sparsity in their latent representations, i.e., different concepts must be able to activate different numbers of latents (Fig. 4).
@@
Figure 4: SpaDE shows adaptive sparsity by projecting onto the probability simplex. In this illustrative $3D$ figure, note $\|\bm{x}\|_{0}=3$ for points on the face, $\|\bm{x}\|_{0}=2$ for points on edges along subspaces, and $\|\bm{x}\|_{0}=1$ for corners on coordinate axes.
@@
To satisfy the desiderata above, we use a simple first-order equality constraint on the projection set $\mathcal{S}$ (Eq. 4), resulting in the probability simplex $\mathcal{S}=\Delta^{s}=\{\bm{x}\in\mathbb{R}^{s}:\sum_{i}x_{i}=1,\bm{x}\geq\bm{0}\}$. The non-negativity is necessary to explain away data as a combination of features with positive contributions. Projection onto the simplex (see Fig. 4) results in the sparsemax nonlinearity ([44]):
$\begin{split}\text{Sparsemax}(\bm{v})&=\operatorname*{arg\,min}_{\bm{\pi}\in\Delta^{s}}\|\bm{\pi}-\bm{v}\|_{2}^{2}.\end{split}$
@@
The probability simplex $\Delta^{s}$ admits representations with any (non-zero) level of sparsity, as illustrated in Fig. 4. Combining Sparsemax with euclidean distances then yields SpaDE:
$\begin{split}\bm{z}=\bm{f}(\bm{x})&=\text{Sparsemax}(-\lambda d(\bm{x},\bm{W})),\\ &\text{where }d(\bm{x},\bm{W}))_{i}=\|\bm{x}-\bm{W}_{i}\|_{2}^{2}.\end{split}$ (5)
@@
In the above, $\lambda$ is a scaling parameter (akin to inverse temperature), while $\bm{W}_{i}$ is the $i^{th}$ column of the encoder matrix $\bm{W}$ which behaves as a prototype (or landmark) in input space since we compute euclidean distance from input $\bm{x}$ to $\bm{W}_{i}$. App. D.4 and D.2.3 describe the receptive fields of SpaDE in further detail and show how it captures nonlinear separability and concept heterogeneity.
@@
We also note that the outer optimization for SpaDE is K-Deep Simplex (KDS, [45]), a modified dictionary learning technique which incorporates locality into sparse representations. The regularizer from KDS is a distance-weighted $\ell_{1}$ regularizer $\mathcal{R}(\bm{z})=\sum_{i}z_{i}\|\bm{x}-\bm{W}_{i}\|_{2}^{2}$, which encourages prototypes to move closer to data when they are active, increasing sparsity of representation  [^2] This regularizer encourages dictionary atoms to “stick” to the data, addressing the recently raised concern [15, 32] that directions learned by SAEs may be out-of-distribution (OOD), contributing to their instability.. The inner optimization for SpaDE is a one-sided sparsity-regularized optimal transport (see App. D.4).
Our claim about the duality between SAE architectures and data assumptions about concepts also applies to SpaDE. Beyond nonlinear separability and concept heterogeneity, SpaDE implicitly assumes that Euclidean distances are useful in concept space—concepts are distance-separated—and distances can be used to disentangle concepts.
@@ 5 Results: Empirical Validation of SAE behavior
We perform a suite of experiments which involve training ReLU, JumpReLU, TopK and SpaDE SAEs on synthetic Gaussian clusters, semi-synthetic formal-language model activations and natural vision model activations. Our synthetic experiments aim to validate our claims about implicit assumptions in SAEs. Experiments on more naturalistic data seek to demonstrate our claims extend to realistic data settings. Further analysis is deferred to App. E. The code to replicate synthetic experiments is available at: https://github.com/Sai-Sumedh/SaeConceptDuality-SpaDE, formal language experiments is at: https://github.com/EkdeepSLubana/spadeFormalGrammars, and vision experiments is at: https://github.com/KempnerInstitute/Overcomplete.
Figure 5: Effect of Nonlinear Separability on SAEs. Each column represents a different SAE. a) $F_{1}$ scores of the top 5 most monosemantic latents (highest F1 scores), where shaded region is $\pm$1SD, of each SAE on two concepts—orange (linearly separable) and purple (non-linearly separable). SAEs that assume linear separability struggle to capture the nonlinearly separable concept. b) Receptive fields of the most monosemantic latent for each SAE, illustrating how some architectures fail to isolate the nonlinear concept cleanly. Intensity of color indicates strength of SAE latent activation. (c) Matrix of pairwise cosine similarities between sparse codes of different datapoints, and data clusters obtained through spectral clustering on this matrix. In the scatter plot, points colored by the same color belong to one spectral cluster, which intuitively indicates that they activate a common set of SAE latents. SpaDE is able to maintain clear concept boundaries and doesn’t mix distinct features, while other SAEs group subsets of different features into the same spectral cluster (same color).
@@ 5.1 Separability Experiment
Dataset and Experiment: We construct a 2-dimensional dataset with Gaussian clusters (abstraction of concepts) of different magnitudes in order to demonstrate nonlinear separability of concepts in a simple setting which facilitates visualization. Here, each cluster is defined as its own concept, and we expect SAEs to learn latents responding to individual clusters. The concepts with smaller norm are not linearly separable, while those with larger norm are linearly separable. We train all SAEs on this dataset for a range of sparsity levels. Following our arguments about implicit assumptions in SAEs, we hypothesize that ReLU and JumpReLU will be unable to capture the nonlinearly separable concepts with monosemantic latents (measured using F1 scores; see Eq. 9).
Observations: Fig. 5 shows how different SAEs fare on this experiment. ReLU and JumpReLU achieve an F1 score of 1 for the separable concept (orange), while their F1 scores are much lower and bounded above (by  0.5) for the nonlinearly separable concept (purple). The receptive fields of ReLU, JumpReLU in Row (b) clearly overlap with other concepts in the nonlinearly separable case. TopK performs somewhat poorly on both concepts, showing comparable F1 scores in both cases. SpaDE shows a top F1 score of 1.0 for both concepts (perfect precision and recall), with its receptive fields capturing concept structure even for nonlinearly separable concepts. While ReLU and JumpReLU show significant cross-concept correlations (between concepts $C_{1},C_{2},C_{3}$, Row (c)) and TopK does marginally better with smaller correlations, SpaDE shows clear delineation of different concepts with clearly separated concepts (no cross correlations, spectral clustering identifies concepts). Note SpaDE may overspecialize and lead to further subclusters, as seen by two colors within concept 1 in row (c).
@@ 5.2 Heterogeneity Experiment
Figure 6: Effect of Concept Heterogeneity on SAEs. a) Per-concept sparsity as a function of intrinsic dimension. Colors indicate per-concept MSE—higher errors (red/yellow) show when an SAE fails to capture a concept effectively. Each solid line indicates one model with a specific choice of hyperparameters. b) Normalized MSE vs. per-concept sparsity. A well-performing SAE should maintain low error across all concepts. TopK SAE only achieves good reconstruction (below the dashed 20% error threshold) when sparsity (fixed for a given model) exceeds intrinsic dimensionality, highlighting its lack of flexibility. c) Cosine similarity between pairs of SAE latents across all concepts (showing co-occurrence), visualized for two sparsity levels.
@@
Dataset and Experiment: We generate Gaussian clusters (again an abstraction for concepts) in a 128-dimensional space. The five concepts are heterogeneous—they belong to subspaces with different intrinsic dimensions (6, 14, 30, 62, 126), but are designed to have isotropic structure within each cluster, and similar total variances across clusters. We trained ReLU, JumpReLU, TopK SAEs and SpaDE on this data with varying sparsity levels. We hypothesize that TopK will not be able to adapt its representations to the intrinsic dimension of each cluster.
@@
Figure 7: Investigating SAE properties on GPT for formal languages. (a) 3D PCA of model activations and SAE encoder weights, where datapoints are colored by part-of-speech (PoS). Encoder weights are indicated by points for SpaDE and arrows for the other SAEs. (b) Matrix of cosine similarities between pairs of data and pairs of latents (in order) for each SAE. White lines separate different PoS. (c) MSE normalized by PoS variance as a function of sparsity, for each PoS. Inset: cumulative sum of variance (eigenvalues of data correlations) of each PoS, where the effective dimension (variance $99\%$) of each PoS is shown. (d) Top-20 $F_{1}$-scores for different PoS from each SAE’s latents (a measure of monosemanticity).
Observations: Fig. 6 shows the results of all SAEs on this experiment. In Row (a), TopK shows the same level of sparsity per concept for all concepts, along with worse reconstruction error for higher dimensional concepts. In contrast, other SAEs—ReLU, JumpReLU and SpaDE—show adaptive sparsity to different extents by adjusting their representations to the intrinsic dimension of each concept. SpaDE can capture the intrinsic dimension nearly perfectly (along the dashed $y=x$ curve) for a specific choice of hyperparameters.
@@
Note that a naïve estimator which predicts the mean of each concept will achieve a normalized MSE of 1. For TopK, normalized MSE (Row (b)) goes below $20\%$ (i.e., explains $80\%$ of the variance) for each concept only when $k$ exceeds the dimension of that concept. For example, $d=6$ goes below the dashed line only after $k=8$, similarly for other concepts. Other SAEs are able to stay below the $20\%$ threshold for nearly all concepts across hyperparameters.
@@
In Row (c), each latent is assigned a concept which it activates maximally for. Note that different concepts use different numbers of latents in ReLU, JumpReLU, and SpaDE. However, there are correlations across concepts in ReLU and JumpReLU (for the dense case), indicating co-occurrence of latents across concepts, which is reduced in the sparse case. Correlations are absent in SpaDE under both cases. TopK uses similar number of latents across concepts, inline with its lack of adaptivity.
@@ 5.3 Formal Languages
Dataset and Experiment: Building on recent work using formal languages for making predictive claims about language models [46, 47, 48], we use this setting as a semi-synthetic setup for corroborating our claims. Specifically, we analyze the English PCFG (Probabilistic Context-Free Grammars, formal models of language often used to study its syntactic properties, see App. C.3) with subject-verb-object sentence order proposed in [34]. We train 2-layer Transformers [49] from scratch on strings of maximum length 128 tokens from the formal grammar above. SAEs are then trained on activations retrieved from the middle residual stream of the model.
Figure 8: SAE properties on DINOv2 activations. (a) 3-D PCA of model activations colored by class, and SAE encoder weights (points for SpaDE, arrows for other SAEs). (b) Cosine similarities of sparse codes of pairs of data and pairs of SAE latents (in order) for each SAE. White lines separate classes. (c) $F_{1}$ scores of top-5 most monosemantic latents for each SAE across classes (color-coded)
Observations: Results are shown in Fig. 7. Different parts of speech (PoS), the core concepts of the grammar, form clusters in a 3D PCA of their representations (see row (a)). SpaDE learns to tile the PoS clusters well. While all SAEs do a good job at making their latents uncorrelated across PoS (first column per SAE, row (b)), there are co-occurring latents across PoS in all SAEs except SpaDE (second column per SAE, row (b)). PoS seem to have different intrinsic dimensions (number of dimensions to capture $99\%$ of total variance in data, inset in row (c)), which leads to TopK requiring different values of K to explain the data (crosses $5\%$ normalized MSE with differing values of k, row (c)). PoS also appear to have differing levels of linear separability, as ReLU and JumpReLU show lower F1 scores which peak at different levels of sparsity for each concept (row (d)), while SpaDE shows a perfect F1 score of 1 in its most monosemantic latents.